OpenStudy (anonymous):

Suppose a landowner wishes to use 3 miles of fencing to enclose an isosceles triangular region of as large an area as possible. What should be the lengths of the sides of the triangle?

OpenStudy (anonymous):

what's the area of an isosceles triangle?

OpenStudy (anonymous):

it doesnt tell

OpenStudy (ryan123345):

Hey there! I see that you are new to openstudy. This is a great website on getting help with questions of any subject. It is a fun way of learning! I wish you good luck on openstudy and dont forget to read the http://openstudy.com/code-of-conduct so you know the rules and regulations :) Hope you have fun!

OpenStudy (anonymous):



OpenStudy (anonymous):

\[h = \sqrt{a ^{2} - \frac{ 1 }{ 4 }b ^{2}}\] by pythagoras

OpenStudy (anonymous):

yes, i got to that

OpenStudy (anonymous):

area will be \[\frac{ 1 }{ 2 }bh\]

OpenStudy (anonymous):

you know what h is, so substitute that

OpenStudy (anonymous):

also you're given from the ques that 2a + b = 3

OpenStudy (anonymous):

from that you can find a in terms of b

OpenStudy (anonymous):

hence area A = \[\frac{ 1 }{ 2 }b \sqrt{\frac{ 3-b }{ 2 } - \frac{ 1 }{ 4 }b ^{2}}\]

OpenStudy (anonymous):

differentiate A w.r.t b and set it equal to 0 (first derivative test)

OpenStudy (anonymous):

is that clear?

OpenStudy (anonymous):

I think finding the right equation was the main thing, after that you can apply the 1st and 2nd derivative rules to find the max or min

OpenStudy (anonymous):

but the problem is i got a lot of trouble when differentiating it

OpenStudy (anonymous):

because the equation is overly complex, and i can never get a resonable answer

OpenStudy (anonymous):

Okay here's what I get after simplifying A \[A = \frac{ 1 }{ 4 }b \sqrt{9-6b}\]

OpenStudy (anonymous):

I think you can easily differentiate this?

OpenStudy (anonymous):

thanks, let me try first

OpenStudy (anonymous):

\[A' = \frac{ -6 }{ 8 }\frac{ b }{\sqrt{9-6b} } + \frac{ 1 }{ 4 }\sqrt{9-6b}\]

OpenStudy (anonymous):

set A' = 0, and find the values of b

OpenStudy (anonymous):

\[\frac{ 1 }{ 2 }b \sqrt{\frac{ 3-b }{ 2 }-\frac{ 1 }{ 4 }b ^{2}}\] are you sure about (3-b)/2 part?